OpenStudy (anonymous):

3(x + 1) = −2(x − 1) + 6

OpenStudy (anonymous):

Distribute. 3(x + 1) = -2(x - 1) + 6 3x + 3 = -2x + 2 + 6 Combine like terms. Solve for x.